OpenStudy (anonymous):

You draw 2 cards from a well-shuffled deck of 52 cards. After you draw the first card, you do not replace it in the deck. The probability of drawing the ace of spades on your first draw is 1/52. What is the probability of P(get queen of hearts on second draw | get ace of spades on first draw)? Write fractions using the slash ( / ) key. Reduce fractions to their lowest terms.

OpenStudy (anonymous):

1/52 x 1/49 = 1/2548? is that the answer?

OpenStudy (anonymous):

Drawing an ace of spade: 1/52 Drawing a queen of hearts: 1/51 (Ace did not get replace) 1/52 * 1/51 =

OpenStudy (anonymous):

oh yeah, i have miss understood the answer a bit

OpenStudy (anonymous):

so the answer would be 1/2652 ?

OpenStudy (anonymous):

yes